A force of 8.0 N is applied for 1.5 s to a model train, having a mass of 1.2 kg, at rest on a horizontal surface. What is the speed of the model train after the 1.5 s time interval?


1
Expert's answer
2021-01-10T18:29:13-0500

We can find the speed of the model train after the 1.5 s time interval from the impulse-momentum change equation:


"F\Delta t=m\Delta v,""F\Delta t=m(v_f-v_i)=mv_f,""v_f=\dfrac{F\Delta t}{m}=\dfrac{8.0\ N\cdot 1.5\ s}{1.2\ kg}=10\ \dfrac{m}{s}."

Answer:

"v_f=10\ \dfrac{m}{s}."